SPECTROSCOPIC INVESTIGATION OF THE LATTICE DYNAMICS IN HYDROGEN-BONDED CRYS-TAL NH_{4}IO_{3}

\underline{ALEXANDER BARABASH}, T. GAVRILKO, G. PUCHKOVSKAYA, and A. ROSHCHIN, Institute of Physics, National Academy of Sciences of Ukraine, 46, Prosp. Nauki, 252022 Kiev, Ukraine; J. BARAN, Institute of Low Temperature and Structure Research of the Polish Academy of Sciences, ul. Okolna, 2, 50 - 950 Wroclaw, Poland.

The lattice dynamics of the ammonium iodate crystal NH_{4}IO_{3} has been investigated by means of IR, Raman and NQR methods near the phase transition temperature 367 K (T_{c}). It is known, that NH_{4}IO_{3} crystal grows from water solution and has orthorhombic syngony (C_{2v}^{9}, z = 4). In contrast with other hydrogen-bonded iodate crystals, which undergo the order-disorder phase transitions, the phase transition in NH_{4}IO_{3} crystal has clear marks of the displacive type phase transition. First, ^{127}I NQR measurements allow to assign the low frequency bands observed in IR and Raman spectra of NH_{4}IO_{3} crystal to the lattice vibrational associated with librational and translational vibrations of the NH_{4}^{+} and IO_{3}^{−} ions (ν < 120 cm^{−1}) and with internal vibrations of N−H and I−O ions (ν < 850 cm^{−1}), whereas high frequency bands were associated with NH...O hydrogen bond. Second, the potential barriers for such vibrations were estimated from the spectroscopic data. Further, it was shown, that the temperature dependence of the asymmetry parameter of the ^{127}I in the pretransition temperature range ((T_{c}- 50) ÷ T_{c}) can not be described by a simple Arrheniuse’s law. In the present paper we attempted to explain this deviation by the decreasing of potential barriers for reorientational vibrations of NH_{4}^{+} and IO_{3}^{−} ions due to the temperature expansion of the crystal volume and the change in the hydrogen bonds network in the pretransition temperature range.